OpenStudy (18jonea):

Determine the radius of the circle. Use the point (-11, 6) on the circle to find the radius. and the center is -5, 2 @jdoe0001

OpenStudy (jdoe0001):

hehehe

OpenStudy (18jonea):

what? @jdoe0001

OpenStudy (jdoe0001):

 notice, the radius is the "distance" between the center and the point "on" the circle thus use the distance formula then \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) &({\color{red}{ -11}}\quad ,&{\color{blue}{ 6}})\quad % (c,d) &({\color{red}{ -5}}\quad ,&{\color{blue}{ 2}}) \end{array}\qquad % distance value d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2}\)

OpenStudy (18jonea):

what is d

OpenStudy (jdoe0001):

yeah... what is it, use the distance formula, "d" is just the "distance" or radius in this case

OpenStudy (18jonea):

-5 +11= 6 2-6 = -4

OpenStudy (18jonea):

@jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (18jonea):

so 36+16 = 52

OpenStudy (18jonea):

@jdoe0001

OpenStudy (jdoe0001):

yeap

OpenStudy (18jonea):

approximately 7.21

OpenStudy (18jonea):

@jdoe0001

OpenStudy (jdoe0001):

well...if you wan to round it up, yes

OpenStudy (18jonea):

ok thanks

OpenStudy (jdoe0001):

yw

OpenStudy (18jonea):

lastly how would you write the standard equation for this

OpenStudy (18jonea):

@jdoe0001

OpenStudy (18jonea):

@Nnesha

OpenStudy (18jonea):

@jdoe0001

Nnesha (nnesha):

\(\bf (x-{\color{brown}{ h}})^2+(y-{\color{blue}{ k}})^2={\color{purple}{ r}}^2 \qquad center\ ({\color{brown}{ h}},{\color{blue}{ k}})\qquad radius={\color{purple}{ r}}\) plug in the given values ~jdoe0001